-3(0.234-x)=2(x+0.974)
Answer by tutorcecilia(2152)

-3(0.234-x)=2(x+0.974)[Use the distributive property]
-0.702+3x=2x+1.948 [Solve for x]
-0.702+0.702+3x=2x+1.948+0.702
3x=2x+2.65
3x-2x=2x-2x+2.65
x=2.65
.
Check by plugging all of the values back into the original equation
-3(0.234-(2.65))=2((2.65)+0.974)
-3(0.234-(2.65))=2((2.65)+0.974)
-3(-2.416)=2(3.624)
7.248=7.248 [checks out]